OpenStudy (anonymous):

A trapezoid has a height of 5 cm and one base that measures 8 cm. If the area of the trapezoid is 28 cm2, find the length of the other base.

OpenStudy (anonymous):

Area of trapezoid = 0.5h (sum of parallel sides)\ so 28 cm^2 = 0.5*5cm( 8 cm + other side) or 28 = 2.5(8 + other side) 28/2.5 = 8 + other side 11.2 - 8 = other side 3.2 = other side